OpenStudy (anonymous):

Which ratio is equal to 5/6? A. 15/16 B. 25/36 C. 30/36 D. 60/70

OpenStudy (inkyvoyd):

Not A or D. Do it yourself.

OpenStudy (anonymous):

When you multiply the nominator and denominator of a fraction by the same value, the result of the division is still the same.

OpenStudy (hoblos):

30/36 = (5*6)/(6*6) = 5/6

OpenStudy (anonymous):

Thank you :)